OpenStudy (anonymous):

5^x=3^(x+1) How do I solve this type of problem?

OpenStudy (anonymous):

Use logarithms. ln(5^x) = ln(3^(x+1)) x ln(5) = (x+1) ln(3) x (ln5-ln3) = ln3 x = ln3/(ln5-ln3)

OpenStudy (anonymous):

Medal?

OpenStudy (anonymous):

where did you get the ln from? wouldn't it be log?

OpenStudy (anonymous):

ln is log to base e.

OpenStudy (anonymous):

i know that, but why did you change log to ln?

OpenStudy (anonymous):

It is written as "ln" too. Same thing.

OpenStudy (anonymous):

LN

OpenStudy (anonymous):

ok, i understand the first part of solving it, but then how do you get to the next one? (3rd line)

OpenStudy (anonymous):

x ln(5) = (x+1) ln(3) x ln(5) = x ln(3) + 1*ln(3) x ln(5) - x ln(3) = ln(3) x (ln(5)-ln(3)) = ln(3)

OpenStudy (anonymous):

what property allows you to separate the x+1 on line 2?

OpenStudy (anonymous):

distributive a(b+c) = ab + ac

OpenStudy (anonymous):

but on logs isn't it logxy=logx+logy? is that the same as distributive?

OpenStudy (anonymous):

In this case, you just have a constant ln(3). Treat it as any constant. And, (x+1) is not "inside the log function". So, treat it as just regular terms.

OpenStudy (anonymous):

ok, so if there were parentheses it wouldn't work?

OpenStudy (anonymous):

If it was 3*ln(x+1), then, you can't write it as 3*ln(x) + 3*ln(1).

OpenStudy (anonymous):

ok, so how do you solve it once you get to x (ln(5)-ln(3)) = ln(3)?

OpenStudy (anonymous):

x = ln3/(ln5-ln3)

OpenStudy (anonymous):

Now, use calculator to find the value.

OpenStudy (anonymous):

ok thanks :)

OpenStudy (anonymous):

2.15 seems to be the answer.